Find the wavelength of the photon emitted when hydrogen makes a transition n = 11 --> 2
Answer #1

from the relation
1/L =R [1/n^2-1/m^2]
1/L = R[1/2^2-1/11^2]
L =484/117 R =3.77x10^-7 m
(R =1.097x10^7 /m)